OpenStudy (anonymous):

in 2,6,10 which term is 106

OpenStudy (saifoo.khan):

find the formula.

OpenStudy (anonymous):

common difference=6-2=4

OpenStudy (saifoo.khan):

4n -2

OpenStudy (saifoo.khan):

106 = 4n -2 108 = 4n n = 27\ therefore: 27th term.

OpenStudy (anonymous):

nth term is given by a+(n-1)d

OpenStudy (saifoo.khan):

v can also do it directly.